A car is taking turn on a banked road to prevent slipping. The road is banked at angle 'θ' with horizontal. Weight of car is acting downwards. Component of weight along the road provides required centripetal force and prevents slipping. Which component of weight will be ' mg cos θ' and which will be mg sin θ?
1
Expert's answer
2020-11-16T07:37:11-0500

Downwards and parrallel to road:


"mg\sin{\theta}"

Normally to the surface of the road:


"mg\cos{\theta}"